OpenStudy (electronia):

y/5 + 10=15? Please help!

OpenStudy (♪chibiterasu):

What are we doing with it.

OpenStudy (electronia):

Solving for y.

OpenStudy (♪chibiterasu):

We have to cancel out terms and isolate y. Do you know how we can cancel out +10?

OpenStudy (electronia):

Wait....Im confused...(im in 7th grade btw)

OpenStudy (♪chibiterasu):

We have to move all terms on the left, except y, to the right so that we are left with y.

OpenStudy (♪chibiterasu):

For example: x + 16 = 30 Solve for x x + 16 - 16 = 30 - 16 x = 14 See what I did there?

OpenStudy (electronia):

OH! I got it. Peiced it together.Thank you. Medal!

OpenStudy (♪chibiterasu):

And remember: multiplication cancels out division. division cancels out multiplication.